Grade 11Magical Mathematics[Interesting Approach]

tantheta=b/athen prove that acos2theta+bsin2theta=a

abhinav
10 Years agoGrade 11

1 Answer

Riddhish Bhalodia
10 Years ago
ok I hope you are familier with the following two identities
cos2\theta = \frac{1-tan^2\theta}{1+tan^2\theta} , sin2\theta = \frac{2tan\theta}{1+tan^2\theta}
now its just the question of using thes two and substituting
tan\theta = b/a
in the equation and you will get the answer very easily